OpenStudy (anonymous):

Which equation is equivalent to 2x + 4a = 10?

OpenStudy (anonymous):

x = 5/4a x = 5 + 2x x = 10 - 4a x = 5 - 2a

OpenStudy (anonymous):

2x + 4a = 10 2x = 10 - 4a x = 5 - 2a

OpenStudy (anonymous):

tyty :D

OpenStudy (anonymous):

no problem